A model corresponding to the cooperative interaction between two species x and y

is given by

dx/dt=(4-2x+y)x

dy/dt=(4+x-2y)y

Find all the equilibrium points of the system and discuss the stability of the system at

these points.


1
Expert's answer
2021-03-05T00:05:10-0500

Equilibrium points occur when both dx/dt = 0 and dy/dt = 0.


"(4-2x+y)x=0"

"(4+x-2y)y=0"

"x=0: 4+0-2y=0=>y=2"

Point "(0, 2)"


"y=0: 4-2x+0=0=>x=2"

Point "(2, 0)"


"\begin{matrix}
 4-2x+y=0 \\
 4+x-2y=0
\end{matrix}"

"x=4, y=4"

Point "(4,4)"


"J=\begin{bmatrix}
 4-4x+y & x \\
 y & 4+x-4y
\end{bmatrix}"

Point "(0, 2)"


"J=\begin{bmatrix}
 6 & 0 \\
 2 & -4
\end{bmatrix}"

Find the eigenvalues


"\begin{vmatrix}
 6-\lambda & 0 \\
 2 & -4-\lambda
\end{vmatrix}"

"( 6-\lambda)( -4-\lambda)=0"

"\lambda_1=6>0, \lambda_2=-4<0"

The equilibrium point "(0, 2)" is unstable.


Point "(2, 0)"

"J=\begin{bmatrix}
 -4 & 2 \\
 0 & 6
\end{bmatrix}"


Find the eigenvalues


"\begin{vmatrix}
 -4-\lambda & 2 \\
 0 & 6-\lambda
\end{vmatrix}"

"( -4-\lambda)(6-\lambda)=0"

"\lambda_1=6>0, \lambda_2=-4<0"

The equilibrium point "(2, 0)" is unstable.


Point "(4, 4)"

"J=\begin{bmatrix}
 -8 & 4 \\
 4 & -8
\end{bmatrix}"


Find the eigenvalues


"\begin{vmatrix}
 -8-\lambda & 4 \\
 4 & -8-\lambda
\end{vmatrix}"

"( -8-\lambda)(-8-\lambda)-16=0"

"\lambda_1=-12<0, \lambda_2=-4<0"

The equilibrium point "(4, 4)" is stable.